OpenStudy (anonymous):

Elena's daughter sold 84 boxes of girl scout cookies. she was told that this amount represented 5% of the troop's sales. How many boxes did the troop sell?

OpenStudy (anonymous):

84=5% of x

OpenStudy (anonymous):

x=8400/5

OpenStudy (anonymous):

5% = 84 100 % = ? 100 * 84 / 5 = 20 * 84 = 1680 boxes of cookies

OpenStudy (anonymous):

oh i perceived is wrongly...

OpenStudy (radar):

Uzma, your perception was correct (same answer)